OpenStudy (anonymous):

y = –3x – 5 The student used –2 as one of the values for the input. Which ordered pair could be taken from the table? (–11, –2) (–2, –11) (–2, 1) (1, –2)

jhonyy9 (jhonyy9):

so than you see ,,as one of the values for the input used (- 2) - in this case write y= -3*(-2) -5 y= 6 -5 y = 1 - this mean (-2,1) - than you see this will be the 3rd choice from these - yes ?